OpenStudy (anonymous):

solve the exponential equation .. express your solutions in exact form only.... 14e xlog9=5 solve the exponential equation .. express your solutions in exact form only.... 14e xlog9=5 @Mathematics

OpenStudy (anonymous):



OpenStudy (anonymous):

got it

OpenStudy (anonymous):

start with \[e^{x\ln(9)}=\frac{5}{14}\] then \[e^{\ln(9^x)}=\frac{5}{14}\] so \[9^x=\frac{5}{14}\]